An X-ray machine emits electromagnetic photons carrying a frequency of 7.5x1017hz. What is is the approximate energy carried by each photon?
1
Expert's answer
2019-08-20T10:05:46-0400

The energy of a photon


"E=h\nu"

"=6.6\times 10^{-34}\:\rm{J}\cdot\rm{s}\times 7.5\times 10^{17}\:\rm{Hz}"

"=5.0\times 10^{-16}\:\rm{J}=3.1\:\rm{keV}"